gray, navy, and tan) and five sizes (extra-small or XS, small or S, medium or M, large or L, extra-large or XL). Larger sizes (e.g., extra-extra-large or XXL) and other colors (e.g., purple) are currently out of stock and must be reordered. Instructions Reconsider the following response data collected from current uniform inventory of unisex scrub tops at Hudson Hospital. The unisex scrub tops in stock include four colors (black, Gray Tan Subtotals XS 4 12 4 2 19 Subtotals Utilize the marginal distribution information as determined in a previous practice assessment. Next, answer the following questions based on this context: one or more of the unisex scrub tops are selected at random to show to new hires. COMMENTS& HINTS: Keep in mind that you will typically be asked to state the answer up to three different ways: first as a fraction in original form, second as a simplified fraction or fraction in reduced form, and third as a percentage with percent symbol, rounded to the nearest whole percentage if necessary. In some cases, you may not be asked to report all three versions of the same answer, so please read carefully.]
Question 19 1 points (Extra Credit) Save Answer Record your response to item #19 here. #19: Given that three unisex scrub tops were selected at random from those that were sized extra-large XL HINTS: This selection must appropriately be made without replacement, agreed? Report this answer as a fraction in original form, stating the fraction horizontally with a slash as the fraction bar. For example, if the answer was 600 out of 800, type only 600/800. what is the likelihood that all three were colored navy? COMMENTS &
Answer #1

total number of extra large =30 and out of them colored navy are 7

hence probability =(7*6*5)/(30*29*28)=210/24360